Suppose in its income statement for the year ended June 30, 2022, The Clorox Company reported the following condensed data (dollars in millions).

Salaries and wages expenses

$ 449

Research and development expense

$ 109

Depreciation expense

90

Income tax expense

275

Sales revenue

5,619

Loss on disposal of plant assets

46

Interest expense

156

Cost of goods sold

3,060

Advertising expense

489

Rent expense

102

Sales returns and allowances

266

Utilities expense

57

Assume the marketing department has presented a plan to increase advertising expenses by $326 million. It expects this plan to result in an increase in both net sales and cost of goods sold of 25%. Assume a tax rate of 20%. (Hint: Increase both sales revenue and sales returns and allowances by 25%.)

a. Prepare a multiple-step income statement. (Round answers to 0 decimal places, e.g. 15,222.)

b. Calculate the gross profit rate and the profit margin. (Round answers to 1 decimal place, e.g. 15.2%.)

Gross profit rate

enter gross profit rate in percentages rounded to 1 decimal place

%

Profit margin

enter profit margin in percentages rounded to 1 decimal place

%
Answer #1
THE CLOROX COMPANY
Multiple Income statement
Sales
Sales Revenue $5,619
Sales Returns And Allowance -$266
Net Sales (5353*125%) $6,691
Cost Of Goods Sold (3060*125%) -$3,825
Gross Profit $2,866
Less: Operating Expenses
Salaries And Wages Expenses $449
Depreciation Expense $90
Advertising Expense (489+326) $815
Research And Development Expense $109
Rent Expense $102
Utilities Expense $57
Total Operating Expense $1,622
Income From Operation $1,244
Other Expense And Losses
Interest Expense $156
Loss On Disposal Of Plants Assets $46
Total Other Expense And Losses $202
Income Before Taxes $1,042
Income Tax Expenses @20% $208
Income After Taxes $834
Part-2
Gross Profit Rate= Gross Profit margin/ Net Sales Revenue
=2866/6691=42.8%
Net Profit Rate= Net Profit margin/ Net Sales Revenue
=834/6691=12.5%